# Cross Layer Optimization and Distributed Reinforcement Learning for Wireless

### v2

@@ I Introduction
360 degree video streaming poses additional challenges to the streaming problem. For instance, the typical size of a 360 degree video chunk is significantly larger than the regular chunk size. Hence, streaming a 360 degree video requires massive amounts of data to be transmitted [3]. Moreover, users typically watch 360 degree videos through virtual reality (VR) head-mounted displays (HMDs). However, the use of HMDs is very sensitive to motion-to-photon (MTP) latency, which needs to be kept below $20$ milliseconds [4]. Reducing the latency calls for video transmission schemes that cut down the amount of data transmitted to only what the user actually sees on her field of view (FoV) [5]. For that, the idea of splitting the 360 degree chunk into tiles and delivering only the tiles that intersect with a user’s FoV has been proposed in the literature [6, 7, 8]. However, wirelessly delivering the FoV portion of the 360 degree chunk in high quality and with tight latency bounds is still very demanding in terms of bandwidth consumption. Therefore, currently existing HMDs resort to either transmitting a low resolution content, or tethering using wires to provide the required high rate. Recently, the use of millimeter wave frequency communications (mmWave) has been proposed as a high data rate wireless medium. MmWave is anticipated to be a major enabler for bandwidth hungry applications such as high quality 360 degree video streaming, as it provides huge spectrum bandwidth.
@@ I-B Contribution
The formulated problem is then decoupled into two subproblems which are solved online per chunk. The first is the physical layer subproblem whose objective is to maximize the download rates for each user while ensuring fairness. The second is the application layer subproblem whose objective is to use the achievable download rate of each user and distribute it among different tiles such that the user’s QoE is maximized. We propose a low complexity algorithm to solve the physical layer subproblem, and show that even though the subproblem is non-convex, the proposed algorithm achieves the optimal solution under certain conditions. On the other hand, due to challenges posed by the application layer subproblem, such as the unavailability of non-causal information (future content) of the user’s FoV, it is hard to find an optimal algorithm to solve it numerically. Therefore, we resort to a distributed RL approach in which an RL agent is considered for each tile. Each RL agent observes an input state and performs an action, while the reward is chosen such that the quality of each tile and the fairness among FoV tiles are both captured (see section V). With this scheme, each RL agent is trained to find the quality decision policy that maximizes the global reward.
@@ V Proposed Algorithm to Solve The Application Layer subproblem
The application layer subproblem is non-convex with non-causal knowledge of the user’s FoV. Moreover, the practical VBR encoding of video tiles makes the quality decisions more difficult. Therefore, we consider a distributed RL approach to find the quality decisions based on real observations. However, considering one RL agent in our problem to make the quality decisions of all tiles will lead to exponential number of actions that the RL needs to choose from for each tile.
To avoid high computation complexity, we propose an approach that uses RL agent for each tile. Therefore, each user will run a number of RL agents that is equal to the number of chunk tiles. At the download of each chunk, the RL agents of user $u$ are run sequentially (one after the other), i.e., RL agent $j$ observes a state $s_{c,j}$ which includes the remaining bandwidth after all tiles $1$ to $j-1$ have decided their download rate (action) and performs an action $a_{c,j}$, where $c$ and $j$ are the chunk and tile indices, respectively. After performing the action, the state transitions to $s_{c+1,j}$, and the RL agent receives a discounted reward. The discounted reward does not only consider the time (as in conventional RL) but also the space (other agents), since the action of the $j$-th RL agent does not only affect the future actions of this agent, but also the actions of all agents $j+1$ to $J$. Moreover, the reward is delayed since it depends on the FoV of the user at the playback of the chunk. The time discounted, space aware, and delayed reward we consider is, $E\{\sum_{c=1}^{C}\sum_{j=1}^{J}\sum_{t=1}^{\infty}\gamma^{t-1}\eta_{c+t}^{j}\}$, where $\eta_{c}^{j}$ is defined in (20). This choice considers the effect of the $j$-th agent’s action on the other tiles and ensures a long term system reward of the current action.
The quality decision of each tile in the next chunk to download is obtained from training a neural network. Each tile’s RL agent observes a set of metrics including the probability that tile $j$ is in the FoV of the chunk to download, as well as the available user’s bandwidth that is obtained by solving the physical layer problem. The agent feeds these parameters to the neural network, which outputs the action. The action is defined by the quality of this tile in the next chunk to download. The total reward is then observed and fed back to the agent to train its neural network model. Our reward function, state, and action spaces are defined as follows:

### v3

@@ I Introduction
360 degree video streaming poses additional challenges to the streaming problem. For instance, the typical size of a 360 degree video chunk is significantly larger than the regular chunk size. Hence, streaming a 360 degree video requires massive amounts of data to be transmitted [3]. Moreover, users typically watch 360 degree videos through virtual reality (VR) head-mounted displays (HMDs). However, the use of HMDs is very sensitive to motion-to-photon (MTP) latency, which needs to be kept below $20$ milliseconds [4]. Reducing the latency calls for video transmission schemes that cut down the amount of data transmitted to only what the user actually sees on her field of view (FoV) [5]. For that, the idea of splitting the 360 degree chunk into tiles and delivering only the tiles that intersect with a user’s FoV has been proposed in the literature [6, 7, 8]. However, wirelessly delivering the FoV portion of the 360 degree chunk in high quality and with tight latency bounds is still very demanding in terms of bandwidth consumption. Therefore, currently existing HMDs resort to either transmitting a low resolution content, or tethering using wires to provide the required high rate. Recently, the use of millimeter wave frequency communications (mmWave) has been proposed as a high data rate wireless medium. MmWave is anticipated to be a major enabler for the fifth-generation (5G) bandwidth hungry applications such as high quality 360 degree video streaming. It promises huge spectrum bandwidth, but comes with implementation predicaments such as sensitivity to signal blockage and beam alignment complexity. Due to the high path loss in such high frequency, beamforming is necessary to form narrow directional beams towards the receiver/transmitter. These narrow beams increase the desired power towards the destination while decreasing the interference levels, at the cost of increasing the mmWave sensitivity to blockage.
@@ I-B Contribution
The formulated problem is then decoupled into two subproblems which are solved online per chunk. The first is the physical layer subproblem whose objective is to maximize the download rates for each user while ensuring fairness. The second is the application layer subproblem whose objective is to use the achievable download rate of each user and distribute it among different tiles such that the user’s QoE is maximized. We propose a low complexity algorithm to solve the physical layer subproblem, and show that even though the subproblem is non-convex, the proposed algorithm achieves the optimal solution under certain conditions. On the other hand, due to challenges posed by the application layer subproblem, such as the unavailability of non-causal information (future content) of the user’s FoV, it is hard to find an optimal algorithm to solve it numerically. Therefore, we resort to a RL based method to undertake the application layer problem. Different from traditional RL approaches, we do not rely on a single RL agent, due to the resulting huge action space and computation overhead. In particular, for a single RL agent per user, the action space will exponentially grow with respect to the number of tiles and quality levels per tile. Alternatively, we consider a distributed RL approach in which an RL agent is considered for each tile. Each RL agent observes an input state and performs an action, while the reward is chosen such that the quality of each tile and the fairness among FoV tiles are both captured (see section V). With this scheme, each RL agent is trained to find the quality decision policy that maximizes the global reward.
@@ V Proposed Algorithm to Solve The Application Layer subproblem
The application layer subproblem is non-convex with non-causal knowledge of the user’s FoV. Moreover, the practical VBR encoding of video tiles makes the quality decisions more difficult. Therefore, we propose an efficient algorithm based on ML. In particular, we consider a distributed reinforcement learning approach to find the quality decisions based on real observations.
Our approach, described in Fig 2, considers a reward that captures the dependency of each agent’s action on the other agents’ actions. Before we describe our model, we first give a brief introduction to RL. In RL, there are two components known as agent and environment in which the agent interacts with the environment, performs some action and receives a reward. The RL observes a state $s_{k}$, and performs an action $a_{k}$. After performing the action, the state of the system transitions to $s_{k+1}$, and the agent receives a reward $R_{k}$. In the learning phase, the agent learns from the rewards it receives at every (state, action) pair $(s_{k},a_{k}$) how good is being at that state $s_{k}$, and what is the best action $a_{k}$ to take. The objective of learning is to maximize the expected cumulative discounted reward defined as $E\{\sum_{k=0}^{\infty}\sum_{t=0}^{\infty}\gamma^{t}\eta_{k+t}\}$.
However, considering one RL agent in our problem to make the quality decisions of all tiles will lead to exponential number of actions that the RL needs to choose from for each tile. To avoid high computation complexity, we propose an approach that uses RL agent for each tile. Therefore, each user will run a number of RL agents that is equal to the number of chunk tiles. At the download of each chunk, the RL agents of user $u$ are run sequentially (one after the other), i.e., RL agent $j$ observes a state $s_{c,j}$ which includes the remaining bandwidth after all tiles $1$ to $j-1$ have decided their download rate (action) and performs an action $a_{c,j}$, where $c$ and $j$ are the chunk and tile indices, respectively. After performing the action, the state transitions to $s_{c+1,j}$, and the RL agent receives a discounted reward. The discounted reward does not only consider the time (as in conventional RL) but also the space (other agents), since the action of the $j$-th RL agent does not only affect the future actions of this agent, but also the actions of all agents $j+1$ to $J$. Moreover, the reward is delayed since it depends on the FoV of the user at the playback of the chunk. The time discounted, space aware, and delayed reward we consider is, $E\{\sum_{c=1}^{C}\sum_{j=1}^{J}\sum_{t=1}^{\infty}\gamma^{t-1}\eta_{c+t}^{j}\}$, where $\eta_{c}^{j}$ is defined in (20). This choice considers the effect of the $j$-th agent’s action on the other tiles and ensures a long term system reward of the current action.
As shown in Fig. 2, the quality decision of each tile in the next chunk to download is obtained from training a neural network. Each tile’s RL agent observes a set of metrics including the probability that tile $j$ is in the FoV of the chunk to download, as well as the available user’s bandwidth that is obtained by solving the physical layer problem. The agent feeds these parameters to the neural network, which outputs the action. The action is defined by the quality of this tile in the next chunk to download. The total reward is then observed and fed back to the agent to train its neural network model. Our reward function, state, and action spaces are defined as follows:
@@
Figure 2: The proposed reinforcement learning based scheduling algorithm for solving the application layer problem.